# Training-Free Sketch-Guided Diffusion with Latent Optimization

### v1

@@ 4.2.2 Real Image Editing
Given a reference sketch, our method can effectively generate new images that follow the layout structure of the source sketch. The visual characteristics and details of the final generated image vary with different initial noises, as shown in Fig. 6. A related study [4] suggests that visual features of an exemplar image can be injected through spatial feature substitution in self-attention blocks. We attempted to introduce their method at every denoising step while optimizing the latent based on cross-attention. That is, by extracting spatial features of the self-attention block obtained from the exemplar image and substituting spatial features in the image generation process, we can ensure that while our method controls the resultant image layout, the visual features of the resulting image are similar to those of the exemplar image. This approach can be applied to real-image editing tasks based on sketches. Fig. 8 demonstrates the results generated by this method. The results indicate that our method can work well with [4] on controlling different elements independently.
@@ 5 Limitations and Discussion
Our method shares the same constraints as Stable Diffusion in generating desired images. Specifically, it encounters primary challenges in the following aspects. First, while our method can synthesize a desired image with a layout similar to the source sketch, there are still some slight differences. As our method heavily relies on the image layout extracted from the cross-attention maps derived from the provided sketch, it may fail if the model cannot accurately recognize the correct layout and object structure, as shown on the left side of Fig. 9. Stable Diffusion is primarily trained on data from natural image domains, limiting its ability to interpret and process sketch images.
This limitation becomes particularly evident when dealing with highly abstract sketches. For instance, for sketches with unenclosed edges, while the human eye can intuitively fill in gaps and imagine the complete form, the model struggles to do so, failing to generate images with the correct layout. In addition, there are cases where the image generation results appear to resemble the provided sketch image, but the depicted object’s structure is misinterpreted. This can lead to the generation of illusionary images, such as the hawk’s misaligned foot positions.
Another limitation we have encountered is handling transformed sketches. In the ImageNet-Sketch dataset, images with different transformations exist. As illustrated on the right side of Fig. 9, our method fails to function effectively if the sketch is flipped from bottom to top. This may be due to the lack of data representing odd angles and positions in the diffusion model training dataset.
@@ 6 Conclusion
We introduce a novel pipeline for the sketch-to-image generation task requiring no model training or fine-tuning. Leveraging the cross-attention mechanism within the T2I diffusion model, we can extract image layout and object structure from the sketch image and utilize extracted core features for new image generation with proposed latent optimization. Our method demonstrates a remarkable balance between preserving the layout structure of the sketch image and aligning with the provided text prompt. Meanwhile, our method can be integrated into other training-free methods, enabling image variation generation and image editing without model fine-tuning. Our work showcases the untapped potential of pre-trained text-to-image models, and we hope it will inspire future research in this direction.

### v2

@@ 4.2.2 Real Image Editing
To further validate performance, we conduct a quantitative evaluation using FID, IoU, and LPIPS on 10k sketches from the ImageNet-Sketch dataset, with ImageNet-1k as the reference for FID calculation. For IoU, we use groundingDINO [15] for object detection. As shown in Table 1, our method achieves the lowest FID (21.54), indicating a high degree of photorealism and better alignment with the natural image distribution.
While our IoU (0.6545) and LPIPS (0.8165) are competitive, we note that these metrics do not fully reflect the perceptual quality in sketch-based generation. For example, although PnP achieves the highest IoU (0.8289) and lowest LPIPS (0.3333), its FID (69.25) is substantially worse—implying over-reliance on sketch structure at the expense of realism.
Given a reference sketch, our method can effectively generate new images that preserve the layout and structure of the input sketch. As shown in Fig. 6, the visual characteristics and fine details of the generated images vary with different initial noise seeds. A related study [4] proposes injecting visual features from an exemplar image via spatial feature substitution in the self-attention blocks. Inspired by this, we integrate their method into our pipeline by performing feature substitution at each denoising step while simultaneously optimizing the latent based on cross-attention maps. Specifically, we extract spatial features from the self-attention blocks of the exemplar image and substitute them into the generation process. This integration allows our method to preserve the structural layout from the sketch, while transferring the visual appearance from the exemplar image. As a result, our pipeline can be extended to support real-image editing tasks guided by sketches. As shown in Fig. 8, the results demonstrate that our method can be effectively combined with [4] to achieve independent control over both structural layout and visual style.
@@ 5 Limitations and Discussion
Our method inherits certain limitations from Stable Diffusion, particularly in generating images that precisely match the desired structure. Although our method is capable of synthesizing images with layouts similar to the input sketch, minor deviations may still occur. This is because our approach relies heavily on the layout information encoded in cross-attention maps derived from the sketch. If the model fails to accurately interpret the sketch’s structure or spatial arrangement, the resulting image may deviate from the intended composition.
This issue is further compounded by the fact that Stable Diffusion is primarily trained on natural image domains, which limits its ability to interpret abstract or stylized sketches effectively. The limitation becomes particularly evident when processing highly abstract sketches, such as those with incomplete or unenclosed edges. While humans can intuitively fill in gaps and recognize the intended object, the model lacks such inference capabilities and may fail to reconstruct the correct layout.
@@ 6 Conclusion
We present a novel training-free pipeline for the sketch-to-image generation task, which requires no model training or fine-tuning. By leveraging the cross-attention mechanism in pre-trained text-to-image diffusion models, our method effectively extracts layout and object structure from a sketch and utilizes these core features for guided image generation through latent optimization.
Our approach achieves a remarkable balance between preserving the spatial structure of the sketch and aligning with the semantic content of the text prompt. Furthermore, it can be integrated with other training-free methods, enabling both image variation generation and real-image editing without model fine-tuning. Our work showcases the untapped potential of pre-trained text-to-image models, and we hope it will inspire future research in this direction.